# Fortune: Formula-Driven Reinforcement Learning for Symbolic Table Reasoning in Language Models

### v2

@@ 3.4 Formula Tuning
Definition. Formula Tuning is a reinforcement learning (RL) framework that defines spreadsheet formulas as an explicit symbolic reasoning space for table understanding. Specifically, we fine-tune a pretrained LM $\pi_{\theta}$ to generate formulas $f\in\mathcal{F}$, which are executed by a deterministic spreadsheet engine $\operatorname{exec}(f,\mathbb{T})$. The resulting answer $a=\operatorname{exec}(f,\mathbb{T})$ is compared against the ground-truth answer $a^{\star}(s)$, and the model receives a reward:
$r(a\mid s)=\begin{cases}1,&\text{if }a=a^{\star}(s),\\[3.00003pt] 0.2,&\text{if }a\neq a^{\star}(s)\text{ and }f\text{ is executable},\\[3.00003pt] 0,&\text{if }f\text{ is not executable}.\end{cases}$ (13)
This reward function encourages the model to explore valid, executable formulas—even if initially incorrect—while assigning full credit only when the answer exactly matches the ground truth.
Objective. Formula Tuning maximizes the expected reward using RL algorithms, such as proximal policy optimization (PPO) [47], with the action space constrained to spreadsheet formulas:
@@
This framework enables the model to perform symbolic reasoning over general tables with higher accuracy, with its advantages analyzed in Section 3.2 and Section 3.3. For additional discussion and practical challenges of the methodology, please refer to Appendix D.
@@ 4.1 Settings
4.2 Performance of Formula Learning under SFT, RL, and Cold-Start RL
Table 1: Performance under Zero-shot, SFT, and RL settings across models and datasets. Values in the table indicate accuracy (%). Text and Formula refer to textual and symbolic reasoning methods, respectively. w/ CS denotes cold-start RL [17] initialized from SFT. For open-source models, the best performance in each column is highlighted in dark blue, and the second-best in light blue.
Base Model Method In-Distribution Out-of-Distribution Overall
WikiTQ TabFact FinQA HiTab MultiHiertt AIT-QA TableBench
Zero-Shot
GPT-4o-mini Text 67.36 88.44 59.20 57.89 22.41 77.67 36.79 58.54
Formula 49.16 74.90 43.07 48.17 28.26 75.53 34.65 50.53
GPT-4o Text 78.57 94.52 63.12 69.26 36.11 81.36 42.66 66.51
Formula 53.36 79.94 48.65 65.40 38.41 85.83 37.92 58.50
O1 Text 77.90 95.50 55.71 74.31 38.31 81.55 45.03 66.90
Formula 70.40 91.11 42.81 75.00 48.95 88.93 44.02 65.89
[…]
Formula 12.58 15.77 22.67 6.57 4.23 20.00 10.31 13.16
[…]
Formula 38.96 52.52 34.44 32.60 15.90 52.43 26.05 36.13
Supervised Fine-Tuning (SFT)
[…]
Formula 59.62 72.48 58.50 72.54 44.00 74.95 31.48 59.08
[…]
Formula 63.46 78.16 58.33 71.83 42.68 76.12 35.56 60.88
Reinforcement Learning (RL)
[…]
Text w/ CS 71.56 87.01 56.84 77.64 49.34 85.66 28.16 65.17
Formula 57.64 80.09 60.85 67.93 29.40 80.78 30.69 58.20
Formula w/ CS 70.49 83.04 71.99 79.29 54.55 81.29 36.64 68.18
[…]
Text w/ CS 71.31 86.07 64.77 77.42 54.25 85.43 25.94 66.46
Formula 67.80 84.19 62.16 71.19 41.72 81.17 35.45 63.38
Formula w/ CS 70.90 86.18 69.21 77.89 56.78 79.14 39.25 68.48
Table 4.2 presents the performance of different formula learning methods under supervised fine-tuning (SFT), reinforcement learning (RL), RL with a cold-start strategy (RL w/ CS), and direct zero-shot inference without any training. This experiment is primarily designed to validate the theoretical analysis discussed earlier and to discuss how its behavior in practical scenarios aligns with our theoretical analysis. Several key analyses and insights are summarized below:
[…]
Formula-tuned small models could surpass closed-source LMs. After SFT + RL, small open-source models achieve strong overall accuracies (68.18% and 68.48%). The performance surpasses GPT series models and even outperforms the large reasoning model O1 with 66.9% accuracy. These results highlight the power of formula tuning in democratizing high-quality table reasoning.
@@ 4.2 Performance of Formula Learning under SFT, RL, and Cold-Start RL
   Binder [36] CodeX 64.60* 85.10* -* -* -*
   Dater [34] CodeX 65.90* 85.60* -* -* -*
   API-Assisted [11] CodeX 42.40* -* 69.30* -* -*
   ReAcTable [67] CodeX 68.00* 86.10* -* -* -*
   Chain-of-Table [35] PaLM 2 67.31* 86.61* -* -* -*
   Norm-DP&Agent [68] GPT-3.5 73.65* 88.50* -* -* -*
   TIDE DP&Agent [69] GPT-3.5 75.00* 89.82* -* -* -*
   TableMaster [11] GPT-4o-mini 78.13* 90.12* -* 66.40* -*
   E5 [70] GPT-4 -* -* 85.08* -* -*
@@
   FORTAP [39] BERT+LSTM -* -* 47.00* -* -*
   TAPEX-${}_{\text{Large}}$ 59.10* 84.20* 45.60* -* -*
@@
4.3 Performance of Fortune and Fortune++ Compared to Other Methods
@@
Table 2 presents the performance of Fortune and Fortune++ compared to several strong baselines, as detailed in Appendix E. Fortune is derived from the best overall performance achieved through cold-start RL in formula-based symbolic reasoning. Following prior work [68, 69, 22], we adopt the self-consistency strategy [54] to enhance table understanding performance. This strategy involves generating multiple candidate formulas and selecting the final answer based on majority voting. It is a widely adopted and effective approach for improving accuracy. Fortune follows previous work by generating 10 symbolic reasoning outputs. To further leverage the complementary strengths of textual and symbolic reasoning, we introduce Fortune++, which produces a balanced mix of 5 textual and 5 symbolic outputs. Neither method relies on a cold-start strategy.
Fortune++ delivers consistently strong performance across benchmarks. Fortune++ surpasses all finetuning-based methods across the reported datasets. Specifically, it achieves 80.47% on FinQA, demonstrating strong complex mathematical reasoning ability. On AIT-QA, Fortune++ brings an improvement of 30.9 points, highlighting the out-of-distribution robustness enabled by RL. These results also show that smaller open-source models can outperform larger closed-source models. Despite using only a 7B-parameter Qwen backbone, Fortune++ consistently outperforms nearly all prompting-based methods, including those powered by GPT-4o. The original Fortune, which relies solely on formula-based reasoning, also achieves competitive performance across benchmarks.
@@
RL surpasses SFT. TabAF [22] is a strong baseline that uses SFT for symbolic reasoning with formula and similarly adopts a hybrid self-consistency strategy with 5 textual and 5 formula-based outputs. Nevertheless, Fortune++ significantly outperforms TabAF, demonstrating that RL offers clear advantages over SFT-only models distilled from stronger teacher models.
@@
Additional results and further analysis. Fortune achieves 40.85% on MultiHiertt and 35.22% on TableBench, while Fortune++ achieves 51.73% and 44.96%, respectively. An ablation study and upper-bound performance analysis of Fortune and Fortune++ are presented in Appendix F. A statistical analysis of the generated formulas is provided in Appendix I, and qualitative case studies are discussed in Appendix J. A comparative study of formula tuning versus other symbolic table reasoning methods is included in Appendix G, and an impact analysis of the reasoning process during formula tuning appears in Appendix H.
@@ 5 Conclusion
In this paper, we introduced Formula Tuning (Fortune), a reinforcement learning framework that trains language models to generate executable spreadsheet formulas for table understanding task. Our findings highlight the promise of formula-driven learning in enhancing reasoning capabilities of language models on tabular tasks. Limitations and future work are discussed in Appendix A.
Appendix A Limitations and Future Work
While Fortune demonstrates strong performance, several limitations remain and suggest promising directions for future research.
Limited datasets and experimental coverage. Due to the resource-intensive nature of reinforcement learning, our evaluation is limited to a few representative public datasets, which primarily consist of clean and well-structured tables. This may not fully capture real-world scenarios, where spreadsheets often contain noisy, irregular, or complex two-dimensional layouts. Additionally, we experimented with only a limited set of base models and reinforcement learning algorithms (e.g., PPO). Nonetheless, we believe the experiments conducted in this work sufficiently demonstrate the effectiveness of our approach. Future work should explore a broader range of model sizes, architectures, and reinforcement learning algorithms across different downstream scenarios, such as applying Formula Tuning to larger models to achieve even better performance.
Applicability to broader table understanding tasks. Our method assumes that answers can be fully derived from tabular data via executable formulas, which holds for many symbolic reasoning tasks. However, this assumption may not extend to tasks involving free-form text, multi-modal inputs, or ambiguous supervision. Nonetheless, formulas may still serve as useful intermediate representations, auxiliary objectives, or reasoning grounding mechanisms in such settings. Investigating how formula tuning can benefit or integrate with these broader tasks is an important direction.
Extensions to other formula-related tasks. Executable formulas are central not only to reasoning but also to related tasks such as formula completion, correction, and refilling. These tasks could benefit from multi-task learning or joint training alongside formula reasoning. Conversely, using these tasks as pre-training objectives may also enhance symbolic reasoning capabilities via formula. Exploring how these tasks can be unified within a single framework could lead to more powerful and general-purpose symbolic table models.
Cold-start challenges in reinforcement learning. For base models with limited symbolic reasoning capabilities and minimal knowledge of spreadsheet formulas, reinforcement learning from scratch can be unstable. In our experiments, we mitigated this by using the same training corpus for both SFT cold-start and RL. However, curating independent and high-quality cold-start corpora and identifying optimal initialization checkpoints for reinforcement learning remain open challenges.Furthermore, reinforcement learning itself is inherently unstable. Developing practical techniques to stabilize training and improve performance remains a critical area for exploration. Enhancing warm-up strategies and training stability could lead to significantly better RL outcomes.
Reward design for formula optimization. Our current reward signal is based solely on binary execution accuracy. While simple and effective, it overlooks important factors such as formula efficiency, token redundancy, and partial credit. Future work can incorporate more fine-grained reward shaping, including length penalties or structure-aware scoring, to improve both learning stability and the quality of generated formulas.
These limitations point to several promising directions for future research: (1) scaling Formula Tuning to diverse domains and tasks, (2) exploring joint learning of symbolic tasks, (3) developing more stable and adaptive reinforcement learning strategies, and (4) advancing reward engineering for structured output generation.
Appendix B Broader Impacts and Safeguards
@@
Broader Impacts. Formula Tuning (Fortune) introduces a reinforcement learning framework that enhances symbolic reasoning for table understanding via spreadsheet formulas. By improving the ability of language models to reason over tabular data with verifiable, executable outputs, our work offers substantial benefits in domains where transparency and precision are essential—such as education, scientific analysis, finance, and public policy. Executable formulas can provide interpretable and auditable reasoning steps, potentially increasing user trust and reliability in AI-generated decisions involving structured data.
@@
Safeguards. To mitigate such risks, we recommend several safeguards for future use of Fortune and similar symbolic reasoning systems. First, generated formulas should undergo verification through deterministic execution engines to ensure correctness. Second, evaluations should be conducted across diverse domains and spreadsheet structures, particularly including noisy or adversarial formats. Third, human-in-the-loop validation should be used in high-stakes applications (e.g., healthcare or financial audits) to ensure interpretability and safety. Finally, we advocate for transparent reporting of formula generation limitations and the inclusion of provenance indicators that show how a particular output was derived, enabling error tracing and accountability.

### v3

@@ 3.2 Formula Tuning
Definition. Formula Tuning is a reinforcement learning (RL) framework that defines spreadsheet formulas as an explicit symbolic reasoning space for table understanding. Specifically, we train a pretrained LM $\pi_{\theta}$ to generate formulas $f\in\mathcal{F}$, which are executed by a deterministic spreadsheet engine $\operatorname{exec}(f,\mathbb{T})$. The reward is defined in two stages: execution success and answer correctness. First, the generated formula must be executable. If execution succeeds, the resulting answer $a=\operatorname{exec}(f,\mathbb{T})$ is compared with the ground-truth answer $a^{\star}(s)$. Formally, the model receives the following reward:
[…]
This reward function encourages the model to explore valid and executable formulas, even when they do not initially produce the correct answer, while assigning full credit only to formulas whose execution results exactly match the ground truth.
Objective. Formula Tuning maximizes the expected reward using RL algorithms, such as proximal policy optimization (PPO) (Schulman et al., 2017), with the action space constrained to spreadsheet formulas:
@@
This framework enables the model to perform symbolic reasoning over complex tables with more accurate numerical computation. Additional methodological details are provided in Appendix D. We further present a theoretical discussion comparing textual and symbolic reasoning in table reasoning, as well as SFT and RL in symbolic table reasoning, in Appendix B.
@@ 4 Experiments
In this section, we train Formula-R1 based on this framework. We present extensive experiments to empirically demonstrate the effectiveness of Formula Tuning, and further provide additional analyses, discussion, and insights based on these observations.
@@ 4.1 Settings
   Binder (Cheng et al., 2023) CodeX 64.60* 85.10* -* -* -*
   Dater (Ye et al., 2023) CodeX 65.90* 85.60* -* -* -*
   API-Assisted (Cao and Liu, 2025) CodeX 42.40* -* 69.30* -* -*
   ReAcTable (Zhang et al., 2023) CodeX 68.00* 86.10* -* -* -*
   Chain-of-Table (Wang et al., 2024) PaLM 2 67.31* 86.61* -* -* -*
   Norm-DP&Agent (Liu et al., 2023) GPT-3.5 73.65* 88.50* -* -* -*
   TIDE DP&Agent (Yang et al., 2025) GPT-3.5 75.00* 89.82* -* -* -*
   TableMaster (Cao and Liu, 2025) GPT-4o-mini 78.13* 90.12* -* 66.40* -*
   E5 (Zhang et al., 2024c) GPT-4 -* -* 85.08* -* -*
@@
   FORTAP (Cheng et al., 2022) BERT+LSTM -* -* 47.00* -* -*
@@
[…]
4.2 Performance of Formula-R1 and Formula-R1-Plus Compared to Other Methods
Table 1 presents the performance of Formula-R1 and Formula-R1-Plus compared to several strong baselines, as detailed in Appendix E. Formula-R1 is derived from the best overall performance achieved with Formula Tuning under the cold-start RL setting.
@@ 4.2 Performance of Formula-R1 and Formula-R1-Plus Compared to Other Methods
For a fair experimental comparison, we adopt the self-consistency strategy (Wang et al., 2023) to enhance table understanding performance following prior work (Liu et al., 2023; Yang et al., 2025; Wang et al., 2025). This strategy generates multiple candidate formulas and selects the final answer through majority voting, and has been widely adopted as an effective approach for improving accuracy in table reasoning. Specifically Formula-R1 follows previous work (Wang et al., 2025) by generating 10 symbolic reasoning outputs. To further leverage the complementary strengths of textual and symbolic reasoning, we introduce Formula-R1-Plus, which produces a balanced set of 5 textual and 5 symbolic outputs. Formula-R1-Plus incorporates both textual and symbolic reasoning components trained with RL, where the textual component is optimized with RL from a cold start using binary answer correctness as the reward signal.
@@
RL surpasses SFT in numerical reasoning over complex tables. TabAF (Wang et al., 2025) is a strong baseline that employs SFT for symbolic reasoning via spreadsheet formulas and similarly adopts a hybrid self-consistency strategy with 5 textual and 5 formula-based outputs. Nevertheless, Formula-R1-Plus significantly outperforms TabAF, demonstrating that RL offers clear advantages over SFT-only models, even when those models are distilled from stronger teacher models. These gains are particularly pronounced on benchmarks that require substantial numerical reasoning and complex table understanding, such as binary verification in TabFact, multi-step numerical computation in FinQA, and in-the-wild complex table reasoning in AIT-QA, where Formula-R1-Plus surpasses the SFT-based TabAF by approximately 12, 35, and 29 points in accuracy, respectively.
Formula-R1-Plus delivers consistently strong performance. Formula-R1-Plus surpasses all finetuning-based methods across the reported datasets. Specifically, it achieves 80.47% on FinQA, demonstrating strong complex mathematical reasoning ability. On AIT-QA, Formula-R1-Plus brings an improvement of 30.9 points, highlighting the out-of-distribution robustness enabled by RL. These results also show that smaller open-source models can outperform larger closed-source models. Despite using only a 7B-parameter Qwen backbone, Formula-R1-Plus consistently outperforms nearly all prompting-based methods, including those powered by GPT-4o. The original Formula-R1, which relies solely on formula-based reasoning, also achieves competitive performance across benchmarks.
Besides, Formula-R1 achieves 40.85% on MultiHiertt and 35.22% on TableBench, while Formula-R1-Plus achieves 51.73% and 44.96%, respectively.
Table 2: Performance under Zero-shot, SFT, and RL settings across models and datasets. Values in the table indicate accuracy (%). Text and Formula refer to textual and symbolic reasoning methods, respectively. The best performance in each column is highlighted in dark blue, and the second-best in light blue.
Base Model Method In‐Distribution Out‐of‐Distribution Overall
WikiTQ TabFact FinQA HiTab MultiHiertt AIT-QA TableBench
Zero‐Shot
[…]
Formula 12.58 15.77 22.67 6.57 4.23 20.00 10.31 13.16
[…]
Formula 38.96 52.52 34.44 32.60 15.90 52.43 26.05 36.13
Supervised Fine‐Tuning (SFT)
[…]
Formula 59.62 72.48 58.50 72.54 44.00 74.95 31.48 59.08
[…]
Formula 63.46 78.16 58.33 71.83 42.68 76.12 35.56 60.88
Reinforcement Learning (RL)
[…]
Formula 57.64 80.09 60.85 67.93 29.40 80.78 30.69 58.20
[…]
Formula 67.80 84.19 62.16 71.19 41.72 81.17 35.45 63.38
RL with Cold-Start SFT (RL w/ CS)
[…]
Formula 70.49 83.04 71.99 79.29 54.55 81.29 36.64 68.18
[…]
Formula 70.90 86.18 69.21 77.89 56.78 79.14 39.25 68.48
4.3 Performance of Formula Learning under SFT and RL
Table 2 presents the performance of different formula learning methods under supervised fine-tuning (SFT), reinforcement learning (RL), reinforcement learning with a cold-start SFT initialization (RL w/ CS), and direct zero-shot inference without any task-specific training. This experiment is primarily designed to compare RL and SFT, as well as textual and symbolic reasoning, in table understanding scenarios. Several key findings and insights are summarized below:
[…]
4.4 Comparative Analysis of Spreadsheet Formulas, SQL, and Python for Symbolic Table Reasoning
In this section, we analyze the use of formulas as a symbolic reasoning tool, and compare them with SQL and Python. Table 3 compare three symbolic reasoning paradigms (SQL queries, Python snippets, and spreadsheet formulas) under both zero-shot and reinforcement learning (RL) settings. To ensure a fair comparison, we train and evaluate these symbolic tools only on datasets with flat, relational tables: WikiTQ, TabFact, and FinQA for training, and all three plus TableBench for evaluation. This restriction is necessary because SQL operates solely on flat tables, while Python/Pandas are also usually worked for flat tabular structures. Several patterns emerge from the comparison:
Method WikiTQ TabFact FinQA TableBench Overall
Zero-Shot
   SQL 17.07 21.15 1.51 7.64 11.84
   Python 30.96 60.39 34.87 16.33 35.64
   Formula 38.96 52.52 34.44 26.05 37.99
Reinforcement Learning (RL)
   SQL 67.58 83.94 38.79 32.09 55.60
   Python 70.46 84.42 65.85 35.26 64.00
   Formula 70.67 84.50 65.89 35.84 64.23
4.5 Additional results and further analysis.
@@ 4.4 Comparative Analysis of Spreadsheet Formulas, SQL, and Python for Symbolic Table Reasoning
A reward sensitivity analysis is presented in Appendix F, where we further examine the choice of 0.2 as the partial reward value in our main experiments. An ablation study and upper-bound performance analysis of Formula Tuning are presented in Appendix G. An impact analysis of the reasoning process during formula tuning appears in Appendix H. A statistical analysis of the generated formulas is provided in Appendix I, and qualitative case studies are discussed in Appendix J.
@@ 5 Conclusion
In this paper, we introduced Formula Tuning (Fortune), an RL framework that trains LLMs to generate executable spreadsheet formulas for table understanding tasks. Based on this framework, we developed Formula-R1 and demonstrated that formula-driven RL can substantially enhance the symbolic reasoning capabilities of LLMs over complex tables with numerical computation. Extensive experiments across multiple benchmarks show that our approach is particularly effective for numerical computation and complex table reasoning, while also yielding strong out-of-distribution generalization. Our findings highlight the promise of formula-driven learning as a practical and effective paradigm for improving LLM reasoning on tabular tasks.
Appendix A Ethical Statement
@@
Formula Tuning (Fortune) introduces a reinforcement learning framework that enhances symbolic reasoning for table understanding via spreadsheet formulas. By improving the ability of language models to reason over tabular data with verifiable, executable outputs, our work offers substantial benefits in domains where transparency and precision are essential—such as education, scientific analysis, finance, and public policy. Executable formulas can provide interpretable and auditable reasoning steps, potentially increasing user trust and reliability in AI-generated decisions involving structured data.
@@
To mitigate such risks, we recommend several safeguards for future use of Fortune and similar symbolic reasoning systems. First, generated formulas should undergo verification through deterministic execution engines to ensure correctness. Second, evaluations should be conducted across diverse domains and spreadsheet structures, particularly including noisy or adversarial formats. Third, human-in-the-loop validation should be used in high-stakes applications (e.g., healthcare or financial audits) to ensure interpretability and safety. Finally, we advocate for transparent reporting of formula generation limitations and the inclusion of provenance indicators that show how a particular output was derived, enabling error tracing and accountability.
Appendix B Theoretical Analysis and Discussion
In this section, we present a theoretical analysis and discussion comparing textual versus symbolic reasoning in table understanding, as well as supervised fine-tuning (SFT) versus reinforcement learning (RL) in symbolic table reasoning (Figure 2).
B.1 Textual vs. Symbolic Reasoning in Table Understanding
Definition 1 (Textual and Symbolic Policies).
  1. 1.
  2. 2.
Figure 2: A simplified illustration contrasting Textual versus Symbolic Reasoning in Table Understanding, and Supervised Fine-Tuning (SFT) versus Reinforcement Learning (RL) in Symbolic Table Reasoning.
Theorem 1 (Symbolic Reasoning Superiority).
Under mild assumptions, the expected reward achieved by symbolic reasoning is greater than or equal to that of textual reasoning for any input $s$:
The assumptions and the proof of Theorem 1 are provided in Appendix C.1.
Remark 1 (Symbolic Reasoning Potential Benefit).
Maximizing the expected reward in Eq. (2) therefore tends to favor the symbolic policy $\pi_{\theta}^{\mathrm{sym}}$ over the textual policy $\pi_{\theta}^{\mathrm{txt}}$. Symbolic reasoning is particularly advantageous for complex tables and questions requiring multi-step computation or precise numerical manipulation, since correctness is determined by the execution result rather than the exact reasoning trace. As a result, it often achieves higher accuracy than purely textual reasoning.
B.2 SFT vs. RL in Symbolic Table Reasoning
Theorem 2 (RL Superiority).
Under mild assumptions, and assuming the reward function $r(a\mid s)$ is reasonably aligned with task success (e.g., exact match), reinforcement learning (RL) can in principle attain higher expected reward than supervised fine-tuning (SFT):
$\mathbb{E}_{s\sim p,\,a\sim\pi_{\theta}^{\mathrm{RL}}}[r(a\mid s)]\;\geq\;\mathbb{E}_{s\sim p,\,a\sim\pi_{\theta^{\star}}^{\mathrm{SFT}}}[r(a\mid s)].$ (6)
The assumptions and the proof of Theorem 2 are provided in Appendix C.2.
Remark 2 (RL Objective and Potential Benefit).
Unlike SFT, which is constrained to imitating the teacher policy $\pi_{g}$, reinforcement learning (RL) directly seeks to maximize the expected task reward:
This objective may allow the model to assign probability mass to high-reward actions that lie outside the support of $\pi_{g}$—for example, alternative formulas that yield the correct answer but differ syntactically or structurally from those observed during supervised training.
In symbolic table reasoning, such flexibility can be particularly helpful: since many distinct formulas can yield the same correct result, RL may leverage this many-to-one mapping by exploring diverse yet semantically valid expressions. Consequently, RL has the potential to surpass the SFT reward bound, under mild assumptions on reward alignment and exploration quality.